How may ounces of a 30% alcohol solution must be mixed with 9 ounces of a 35% alcohol solution to make a 31% alcohol solution?
Answer by stanbon(75887)

How may ounces of a 30% alcohol solution must be mixed with 9 ounces of a 35% alcohol solution to make a 31% alcohol solution?
------
alcohol + alcohol = alcohol
0.30x + 0.35*9 = 0.31(x+9)
----
Multiply thru by 100 to get:
30x + 35*9 = 31x + 31*9
-x = -4*9
x = 36 ounces (amt. of 30% solution needed)
=================
Cheers,
Stan H.